OpenStudy (anonymous):

2 part time jobs. First job is $8hr at library. second job is $15hr for construction. Works 25 hrs per week. Enjoy's working at library more. How many hours can he work at library at earn at least $300 week?

OpenStudy (jamesj):

Let L be the number of hours working at the library and C the number in construction. First note that L + C = 25, so C = 25 - L Now the total amount earned is 8L + 25C = 8L + 25(25 - L) = 575 - 17L and we want that that amount to be at least 300 So 575 - 17L >= 300 275 >= 17 L 17L =< 275 L =< 4.41 So he should work at the library no more than ~4.4 hours.

OpenStudy (anonymous):

do you factor the rate the student makes in construction @$15?

OpenStudy (jamesj):

Oh whoops, I thought it was $25. Well, same calculation, slightly different answer.

OpenStudy (jamesj):

If you use $15 for Construction job, you end up with 7L < 75 So L < 10.7. Hence no more than 10.7 jobs. Sorry about that.

OpenStudy (anonymous):

cheers. :)